How many significant figures are in the following measured quantity? express the answer as an integer.

0.0003303 m
1
Expert's answer
2020-12-16T04:53:32-0500

There are 4 significant figures in 0.0003303 m.

0.0003303 m = 3303 × 10^-7 m

Recall that:

100 nanometer (nm) = 10^-7 m

x nm = 3303 × 10^ -7 m

(cross multiply)

x nm = (100 nm × 3303 × 10^ -7 m) / (10^-7 m)

x nm = 0.03303 / 10 ^-7

x nm = 330300 nm.

Hence, the integer value for 0.0003303 m is 330300 nm.